(7,-1);9x-4y=7
Write an equation of the line containing the given point and parallel to the given line.

Answer by rfer(16322)

9x-4y=7
-4y=-9x+7
y=9/4x-7/4
y=2.25x-1.75
-----------
-1=2.25*7+b
-1=15.75+b
b=-16.75
------------
y=2.5x-16.75